If an earthquake measured 5.9 on the richter scale, what was the intensity of the earthquake? Neeed to use the formula M=log(A/Ao).: If an earthquake measured 5.9 on the richter scale, what was the intensity of the earthquake? Neeed to use the formula M=log(A/Ao).
Answer by stanbon(19682)

If an earthquake measured 5.9 on the richter scale, what was the intensity of the earthquake? Neeed to use the formula M=log(A/Ao).
------------------------
5.9 = log(A/Ao)
A/Ao = 10^5.9
A = 10^5.9 Ao
----------------
Intensity was 10^5.9 times the minimum measureable intensity of the sensing
equipment.
------------------
Cheers,
Stan H.